# ViMoE: An Empirical Study of Designing Vision Mixture-of-Experts

### v1

@@ Abstract
Mixture-of-Experts (MoE) models embody the divide-and-conquer concept and are a promising approach for increasing model capacity, demonstrating excellent scalability across multiple domains. In this paper, we integrate the MoE structure into the classic Vision Transformer (ViT), naming it ViMoE, and explore the potential of applying MoE to vision through a comprehensive study on image classification. However, we observe that the performance is sensitive to the configuration of MoE layers, making it challenging to obtain optimal results without careful design. The underlying cause is that inappropriate MoE layers lead to unreliable routing and hinder experts from effectively acquiring helpful knowledge. To address this, we introduce a shared expert to learn and capture common information, serving as an effective way to construct stable ViMoE. Furthermore, we demonstrate how to analyze expert routing behavior, revealing which MoE layers are capable of specializing in handling specific information and which are not. This provides guidance for retaining the critical layers while removing redundancies, thereby advancing ViMoE to be more efficient without sacrificing accuracy. We aspire for this work to offer new insights into the design of vision MoE models and provide valuable empirical guidance for future research.
@@ 1 Introduction
Figure 1: Top-1 accuracy on ImageNet-1K. We compare ViMoE with other ViT architecture baselines. All models are evaluated at $224\times 224$.
@@
General artificial intelligence is continuously developing toward larger and stronger models (Achiam et al., 2023; Reid et al., 2024; Yang et al., 2024; Dubey et al., 2024). However, larger models require significant computational resources for training and deployment, and balancing performance with efficiency remains a critical issue, especially in resource-constrained environments. A promising approach is to use the Mixture-of-Experts (MoE) (Jacobs et al., 1991; Eigen et al., 2013) layers in neural networks, which decouple model size from inference efficiency. MoE embodies the divide-and-conquer principle, where feature embeddings are routed to selected experts through a gating mechanism, allowing each expert to specialize in a subsets of the data. As a result, each input is processed by only a small portion of the parameters, whereas traditional dense models activate all parameters for every input. This approach is becoming increasingly popular in Natural Language Processing (NLP), as it enables parameter scaling while keeping computational costs at a modest level (Jiang et al., 2024; Dai et al., 2024).
@@
This work focuses on exploring the simple application of MoE in vision models. We convert the classic Vision Transformer (ViT) (Dosovitskiy, 2020) into a sparse MoE structure, naming it ViMoE. Our modification of ViT follows (Riquelme et al., 2021), where the Feed-Forward Networks (FFNs) in each block is replaced with multiple experts, while keeping the structure of each expert the same. For simplicity and efficiency, we choose to select experts at the image level rather than the token level (Daxberger et al., 2023; Liu et al., 2024). Through a comprehensive study on image classification, we explore strategies for configuring MoE in a stable and efficient manner, while also observing several interesting phenomena related to expert routing from different perspectives.
@@
An essential consideration in designing ViMoE is determining how many MoE layers to include and where to position them. A common approach is to insert them into the last $L$ ViT blocks (Wu et al., 2022; Liu et al., 2024), which receive the largest gradient magnitudes. Alternatively, one more straightforward approach would be to add MoE layers to all blocks without careful design. We adopt an exhaustive way of scanning the number of layers to determine which configuration yields the optimal accuracy for ViMoE. Interestingly, increasing the number of MoE layers does not always lead to better performance; instead, a downward trend emerges beyond a certain number of layers. We attribute this to the fact that inappropriate MoE layers, particularly in the shallow ViT blocks, not only fail to contribute but also complicate optimization. While scanning and observing can reveal the optimal performance point and the most suitable number of MoE layers, such an approach is invariably laborious. Inspired by Xue et al. (2022); Dai et al. (2024), we introduce a shared expert that absorbs knowledge from the entire dataset, alleviating the inadequacies in individual expert learning and the burden on the routing mechanism. The shared expert brings more excellent stability to ViMoE, as it prevents the accuracy degradation observed with an excessive number of MoE layers. This eliminates the need for constant trial and error to find the optimal point, thereby facilitating a more streamlined design process.
@@
The above are deductions drawn from the scanning results, but we seek further heuristic exploration. Building on the stable ViMoE, we attempt to delve deeper into the routing behavior within MoE layers to uncover what each expert focuses on. Owing to our routing strategy, we can observe how data from each class are distributed across the experts. For the MoE layers in the deeper ViT blocks, the gating network effectively allocates samples of the same class to the same expert, with each expert specializing in processing different data. However, in the shallow blocks, the gating network struggles to consistently route images of the same class to the same expert or effectively guide the experts to specialize in different classes. This suggests that the experts have not learned highly discriminative knowledge; rather, they end up implementing very similar functions, indiscriminately extracting common features across all classes (Riquelme et al., 2021). These results highlights which layers truly fulfill the divide-and-conquer role and which do not, corresponding to the accuracy trends observed through layer scanning.
@@
Furthermore, we aim to inform more thoughtful and efficient ViMoE designs through our observations of MoE behavior. One attempt we propose is to estimate the necessary number of MoE layers based on the routing distribution, and then combine this with the number of experts set per layer to approximate the required expert combinations. This insight allows us to simplify the structure by removing potentially redundant MoE layers, thereby achieving a more efficient ViMoE. As a result, our ViMoE based on ViT-S/14 outperforms DINOv2 (Oquab et al., 2023) by 1.1% on ImageNet-1K (Deng et al., 2009) fine-tuning. With less than one-third of the activated parameters, ViMoE even surpasses a number of advanced ViT-B/16 models (Touvron et al., 2021; Bao et al., 2021; Zhou et al., 2021; Zhang et al., 2022; Xinlei et al., 2021; He et al., 2022).
@@
In summary, we believe that as MoE becomes more widely adopted in vision tasks, the observations, evidence, and analyses presented in this study are worth knowing. We hope that our insights and experiences will contribute to advancing this frontier.
2 ViMoE
@@ 2.1 Preliminary
Mixture-of-Experts (MoE) (Jacobs et al., 1991; Jordan & Jacobs, 1994) is a promising approach that allows for scaling the number of parameters without increasing computational overhead. For Transformer-based MoE models, the architecture mainly consists of two key components: (1) Sparse MoE Layer: A MoE layer contains $N$ experts (denoted as $E_{i}(\cdot),i=1,2,\ldots,N$), each functioning as an independent neural network (Shazeer et al., 2017). (2) Gating Network: This component is responsible for routing the input token $\bm{x}$ to the most appropriate top-$k$ experts (Cao et al., 2023). The gate consists of a learnable linear layer, defined as $g(\bm{x})=\sigma(\bm{W}\bm{x})$, where $\bm{W}$ is the gate parameter, and $\sigma$ is the softmax function. Let $\mathcal{T}$ represent the set of the top-$k$ indices, and output of the layer is then computed as a linear combination of the outputs from the selected experts weighted by the corresponding gate values,
$\bm{y}=\sum_{i\in\mathcal{T}}g_{i}(\bm{x})\cdot E_{i}(\bm{x}).$ (1)
Load Balancing Loss. To encourage load balancing among the experts, we incorporate a differentiable load balancing loss (Lepikhin et al., 2020; Zoph et al., 2022) into each MoE layer, promoting a more balanced distribution of input tokens across the experts. For a batch $\mathcal{B}$ containing $T$ tokens, the auxiliary loss is calculated as a scaled dot product between the vectors $f$ and $P$,
@@
$f_{i}=\frac{1}{T}\sum_{\bm{x}\in\mathcal{B}}\bm{1}\{\text{argmax }g(\bm{x})=i\},$ (3)
$P_{i}=\frac{1}{T}\sum_{\bm{x}\in\mathcal{B}}g_{i}(\bm{x}).$ (4)
@@
MoE Transformer. A widely used approach to applying MoE to Transformer models is to replace the Feed-Forward Networks (FFNs) in some of the standard (non-MoE) Transformer blocks with MoE layers (Fedus et al., 2022). Specifically, in an MoE layer, the experts retain the same structure as the original FFNs. The gating function receives the output from the preceding self-attention layer and routes the token representations to different experts.
2.2 Settings for Image Classification
@@
Architecture. We introduce a ViMoE framework to facilitate our study on the application of MoE for image classification. We choose the Vision Transformer (ViT) (Dosovitskiy, 2020) backbone and replace the FFNs in the ViT blocks with MoE layers. Instead of training from scratch (Riquelme et al., 2021), we consider inheriting self-supervised pre-training weights, which reduces training costs while also benefiting from advanced feature representations. Since the experts in the MoE layers share the same structure as the FFNs, we simply replicate the pre-trained weights of the FFNs across each expert for initialization.
Routing Strategy. Recent large-scale sparse MoE models (Achiam et al., 2023; Jiang et al., 2024; Dai et al., 2024; Yang et al., 2024) typically employ a token-based routing strategy, where the gating mechanism assigns each token to selected experts. However, it is worth considering whether this strategy is necessary for MoE in image classification, where the model focuses more on the overall features of the image to predict a single class for the image. We suggest that the routing strategy should align with the specific requirements of the vision task. Routing at the image level (i.e., selecting experts for each entire image) (Daxberger et al., 2023; Liu et al., 2024) is simpler and better suited to the objectives of image classification. In practice, we use the [CLS] token to represent the image $\bm{x}$ as the input to the gating network, since it encapsulates the information from all image tokens and is used for classification prediction. Additionally, unless otherwise specified, we default to selecting only the top-1 routed expert to simplify the architecture. Therefore, compared to token-based routing, this strategy reduces the number of experts activated per image.
@@
Shared Expert. There is often some common sense or shared information across input tokens assigned to different experts. As a result, with a conventional routing strategy, multiple experts may acquire overlapping knowledge within their respective parameters. By designing the shared expert (Xue et al., 2022; Dai et al., 2024) to focus on capturing and consolidating common information, other routed experts can specialize in learning unique knowledge, leading to a more parameter-efficient model composed of a greater number of specialized experts. Consequently, we introduce the shared expert $E_{s}(\cdot)$ into ViMoE to enable learning common knowledge from all data. In our implementation, we set the number of shared experts to 1, with a structure identical to that of the other experts. The output of the shared expert is added to the output of the selected routed expert, allowing Eq. 1 to be rewritten as,
$\bm{y}=E_{s}(\bm{x})+\sum_{i\in\mathcal{T}}g_{i}(\bm{x}_{\texttt{[CLS]}})\cdot E_{i}(\bm{x}).$ (5)
@@ 3.1 A Stability Strategy for Convenient Design
Scanning the Number of MoE Layers. When designing ViMoE, an important consideration is determining how many MoE layers to include and where to place them within the ViT blocks. Here, we start by exploring sparse MoE without the shared expert for simplicity. The most straightforward approach is to place the MoE layer in every ViT block or to select the last $L$ blocks where the gradient magnitudes are the largest. To explore reasonable configurations and seek guiding insights, we scan the number of MoE layers and evaluate the accuracy of image classification. Our experiments are based on the DINOv2 (Oquab et al., 2023) pre-trained ViT-S/14 (Dosovitskiy, 2020), modified into ViMoE and fine-tuned on ImageNet-1K (Deng et al., 2009) for $200$ epochs (more implementation details are provided in Sec. 4.1). From Fig. 2, it can be observed that regardless of the number of experts, whether $N=2$, $N=4$, or $N=8$, the accuracy consistently exhibits a trend of initially increasing and then decreasing, with this trend becoming more pronounced as $N$ increases. This phenomenon has also been mentioned in Daxberger et al. (2023). We hypothesize that introducing multiple experts too early in the shallow ViT blocks leads to optimization difficulties, and the gating network struggles to achieve precise routing due to limited information (a more detailed analyze of this is given in Fig. 5). This suggests a potential instability in the design of ViMoE. Simply adding MoE layers to all ViT blocks without careful consideration may not lead to optimal results. A scan over different values of $L$ is required to determine the most suitable number of layers, which inevitably increases the design cost.
Shared Expert for Stabilising ViMoE. As previously discussed, the shared expert learns and consolidates knowledge from all the data, making it more effective in capturing common information. We consider this structure effective in alleviating challenges of gating decisions and the limitations of individual expert learning in sparse structures. Therefore, we attempt to incorporate the shared expert into ViMoE to mitigate the potential instability in training MoE layers. In Fig. 3 we present a comparison between models with and without shared expert. Incorporating the shared expert allows ViMoE to achieve stable results, eliminating the need for an exhaustive search to determine the optimal number of layers $L$. Even the naive approach of adding MoE layers to all ViT blocks yields good accuracy, preventing performance degradation caused by inappropriate MoE configurations. Additionally, with the inclusion of shared expert, ViMoE achieves a 0.4% improvement in accuracy (84.3% vs. 83.9%), and a 1.2% increase compared to the DINOv2 baseline (83.1%).
Figure 3: Top-1 accuracy of ViMoE with and without the shared expert at different values of $L$.
Figure 4: Convergence curves for training ViMoE under different configurations.
Convergence Advantage. Taking $N=8$ and $L=12$ as an example, Fig. 4 shows the training curves with and without shared expert, along with the DINOv2 baseline for reference. It is evident that simply adding sparse MoE layers slows down convergence in the early training epochs, and the final performance is nearly indistinguishable from the baseline, supporting the hypothesis that an improper MoE setting can even hinder optimization. In contrast, when shared expert is introduced, training becomes more stable, convergence is faster, and accuracy improves significantly. It is worth mentioning that, with the introduction of shared expert, the MoE layers contain a total of 9 experts (1 shared expert and 8 routed experts), and the forward pass activates both the shared expert and one selected routed expert. To ensure a fairer comparison, we conducted an ablation study by selecting the top-2 experts from the 9 routed experts. On one hand, selecting 2 out of 9 can be seen as a denser setup compared to selecting 1 out of 8, which partially mitigates the negative effects of being overly sparse. On the other hand, even with the same number of experts and activated experts, shared expert still demonstrates the advantage with faster convergence and higher accuracy.
3.2 Efficient Exploration Based on Stability
After constructing the stable ViMoE, we further analyze Fig. 3 and observe the presence of a performance plateau. Interestingly, the turning point differs for each $N$. For $N=2$, $N=4$, and $N=8$, accuracy already surpasses 84.2% at $L=5$, $L=3$, and $L=2$, respectively. Beyond these number of layers, no significant improvement is observed by adding more MoE. We attempt to explain these phenomena and propose strategies for designing a more efficient ViMoE.
Routing Heatmap. Taking $N=8$ as an example, we plot the routing heatmaps of several MoE layers in Fig. 5. These heatmaps illustrate the distribution of class samples across different experts, helping us observe whether the experts are capable of capturing distinctive information. It can be observed that for the MoE layers in the shallow ViT blocks (e.g., $l=12$), the gating network struggles to consistently route images of the same class to the same expert or effectively distinguish the classes each expert should focus on. This indicates that the experts fail to learn highly discriminative knowledge; instead, they are likely performing similar functions, indiscriminately extracting common features. We then focus on the layer where the accuracy plateau occurs for $N=8$, corresponding to $L=2$. It is evident that in the last two MoE layers, the gating network can effectively assign the appropriate expert to each class, and the multiple experts can specialize in handling the corresponding data. Therefore, we conclude that the deep layers are where MoE truly achieves its divide-and-conquer objective, with different experts specializing in handling class-specific content. This observation validates the empirical approach of placing MoE layers in the last few ViT blocks (Wu et al., 2022; Liu et al., 2024) as a reasonable strategy. In contrast, MoE struggles to demonstrate its advantages in the shallow ViT blocks, as the use of multiple experts seems unnecessary for capturing basic visual features. The sparse structure may instead introduce optimization difficulties, making the original dense FFN structure a simpler and more suitable choice.
@@ 3.2 Efficient Exploration Based on Stability
Figure 5: Routing heatmap for the $l$-th MoE layer, where $l=1$ represents the deepest (last) layer and $l=12$ denotes the shallowest (first) layer. The $x$-axis is the class ID from ImageNet-1K, and the $y$-axis is the expert ID. The label order in each figure is adjusted for better readability. Darker colors indicate a higher proportion of images from the corresponding class routed to the expert.

### v2

@@ Abstract
Mixture-of-Experts (MoE) models embody the divide-and-conquer concept and are a promising approach for increasing model capacity, demonstrating excellent scalability across multiple domains. In this paper, we integrate the MoE structure into the classic Vision Transformer (ViT), naming it ViMoE, and explore the potential of applying MoE to vision through a comprehensive study on image classification and semantic segmentation. However, we observe that the performance is sensitive to the configuration of MoE layers, making it challenging to obtain optimal results without careful design. The underlying cause is that inappropriate MoE layers lead to unreliable routing and hinder experts from effectively acquiring helpful information. To address this, we introduce a shared expert to learn and capture common knowledge, serving as an effective way to construct stable ViMoE. Furthermore, we demonstrate how to analyze expert routing behavior, revealing which MoE layers are capable of specializing in handling specific information and which are not. This provides guidance for retaining the critical layers while removing redundancies, thereby advancing ViMoE to be more efficient without sacrificing accuracy. We aspire for this work to offer new insights into the design of vision MoE models and provide valuable empirical guidance for future research.
@@ 1 Introduction
Figure 1: Top-1 accuracy on ImageNet-1K. We compare ViMoE with other ViT architecture baselines. All models are evaluated at resolution $224\times 224$.
@@
Artificial general intelligence is continuously developing toward larger and stronger models [1, 35, 48, 11]. However, larger models require significant computational resources for training and deployment, and balancing performance with efficiency remains a critical issue, especially in resource-constrained environments. A promising approach is to use the Mixture-of-Experts (MoE) [21, 12] layers in neural networks, which decouple model size from inference efficiency. MoE embodies the divide-and-conquer principle, where feature embeddings are routed to selected experts through a gating mechanism, allowing each expert to specialize in a subset of the data. As a result, each input is processed by only a small portion of the parameters, whereas traditional dense models activate all parameters for every input. This approach is becoming increasingly popular in natural language processing (NLP), as it enables parameter scaling while keeping computational costs at a modest level [22, 6, 47, 28, 31, 27].
@@
This work focuses on exploring the simple application of MoE in vision models. We convert the classic Vision Transformer (ViT) [9] into a sparse MoE structure, naming it ViMoE. Our modification of ViT follows Riquelme et al. [36], where the feed-forward network (FFN) in each block is replaced with multiple experts while keeping the structure of each expert the same. For simplicity and efficiency, we choose to select experts at the image level [7, 29] rather than the token level [36, 33]. Through a comprehensive study on image classification and semantic segmentation, we explore strategies for configuring MoE in a stable and efficient manner, while also observing several interesting phenomena related to expert routing from different perspectives.
@@
An essential consideration in designing ViMoE is determining how many MoE layers to include and where to position them. A common approach is to insert them into the last $L$ ViT blocks [43, 29], which receive the largest gradient magnitudes. Alternatively, one more straightforward approach would be to add MoE layers to all blocks without careful design. We adopt an exhaustive way of scanning the number of layers to determine which configuration yields the optimal classification accuracy for ViMoE. Interestingly, increasing the number of MoE layers does not always lead to better performance; instead, a downward trend emerges beyond a certain number of layers. We attribute this to the fact that inappropriate MoE layers, particularly in the shallow ViT blocks, not only fail to contribute but also complicate optimization. While scanning and observing can reveal the optimal performance point and the most suitable number of MoE layers, such an approach is invariably laborious. Inspired by Xue et al. [46], Dai et al. [6], we introduce a shared expert to absorb knowledge from the entire dataset, alleviating the inadequacies in individual expert learning and the burden on the routing mechanism. The shared expert brings more excellent stability to ViMoE, as it prevents the accuracy degradation observed with an excessive number of MoE layers. This eliminates the need for constant trial and error to find the optimal point, thereby facilitating a more streamlined design process.
@@
The above are deductions drawn from the scanning results, but we seek further heuristic exploration. Building on the stable ViMoE, we attempt to delve deeper into the routing behavior within MoE layers to uncover what each expert focuses on. Owing to our routing strategy, we can observe how data from each class are distributed across the experts. For the MoE layers in the deeper ViT blocks, the gating network effectively allocates samples of the same class to the same expert, with each expert specializing in processing different data. However, in the shallow blocks, the gating network struggles to consistently route images of the same class to the same expert or effectively guide the experts to specialize in different classes. This suggests that the experts have not learned highly discriminative knowledge; rather, they end up implementing very similar functions, indiscriminately extracting common features across all classes [36]. These results highlight which layers truly fulfill the divide-and-conquer role and which do not, corresponding to the accuracy trends observed through layer scanning.
@@
Furthermore, we aim to inform more thoughtful and efficient ViMoE designs through our observations of MoE behavior. One attempt we propose is to estimate the necessary number of MoE layers based on the routing distribution, and then combine this with the number of experts set per layer to approximate the required expert combinations. This insight allows us to simplify the structure by removing potentially redundant MoE layers, thereby achieving a more efficient ViMoE. As a result, our ViMoE based on ViT-S/14 [9] outperforms DINOv2 [32] by 1.1% on ImageNet-1K [8] fine-tuning. ViMoE achieves performance comparable to larger models [39, 40, 2, 51, 49, 45, 19] at a smaller scale, as illustrated in Fig. 1. Furthermore, we validate these observations and conclusions on the semantic segmentation task, confirming their generalizability and broad applicability.
@@
In summary, we believe that as MoE applications in vision tasks expand, the observations, evidence, and analyses in this study are worth knowing. We hope that our insights and experiences will contribute to advancing this frontier.
2 Related Work
Mixture-of-Experts (MoE) [21] has been widely studied for its ability to modularize learning and reduce interference across data domains [26, 52, 34, 16, 53]. MoE uses a gating network to assign which experts should handle each data sample. Early MoE models were densely activated, which was effective but computationally expensive [30]. Modern MoE models [20, 18] can be regarded as an application of dynamic neural networks [17], using sparse activation selecting only a subset of experts per input, which greatly reduces computational costs while maintaining performance. This efficient approach is crucial in NLP, as shown in works like Switch Transformers [14], GShard [25], and GLaM [10], which apply sparse MoE to handle large tasks while optimizing resources.
MoE in Vision Tasks. The efficiency of MoE in NLP has inspired its use in the visual domain. Works such as V-MoE [36] and M3vit [13] integrate sparse MoE architectures into ViT, replacing dense feedforward layers with sparse MoE layers to boost efficiency and performance in image classification. Simultaneously, pMoE [5] and DiT-MoE [15] introduce sparse computation: pMoE uses CNN experts for selective image patch processing, while DiT-MoE enhances input-dependent sparsity in diffusion transformers for better image generation. Additionally, some works [4, 43] focus on multi-task visual recognition and efficient training of large MoE vision transformers.
Transformer for Vision. Transformers first saw great success in NLP and were later adapted for computer vision with Vision Transformers (ViT) [9], which process images as patches (like words in text) for global feature extraction. Unlike convolutional neural networks (CNNs) that rely on local receptive fields, the ViT architecture captures broader context, often matching or surpassing CNN performance. In self-supervised learning, models like MoCov3 [45] adapted momentum contrast to ViT, training high-quality visual features from unlabeled data. Inspired by masked language modeling [24], methods such as BEiT [2], MAE [19], and iBOT [51] use masked image modeling to improve generalization. DINOv2 [32] further advanced self-supervised ViT through knowledge distillation on large datasets.
3 Vision Mixture-of-Experts
@@ 3.1 Preliminary
Mixture-of-Experts (MoE) [21, 23] is a promising approach that allows for scaling the number of parameters without increasing computational overhead. For Transformer-based MoE models, the architecture mainly consists of two key components: (1) Sparse MoE Layer: A MoE layer contains $N$ experts (denoted as $E_{i}(\cdot),i=1,2,\ldots,N$), each functioning as an independent neural network [37]. (2) Gating Network: This component is responsible for routing the input token $\boldsymbol{x}$ to the most appropriate top-$k$ experts [3]. The gate consists of a learnable linear layer, defined as $g(\boldsymbol{x})=\sigma(\boldsymbol{W}\boldsymbol{x})$, where $\boldsymbol{W}$ is the gate parameter, and $\sigma$ is the softmax function. Let $\mathcal{T}$ represent the set of the top-$k$ indices, and output of the layer is then computed as a linear combination of the outputs from the selected experts weighted by the corresponding gate values,
$\boldsymbol{y}=\sum_{i\in\mathcal{T}}g_{i}(\boldsymbol{x})\cdot E_{i}(\boldsymbol{x}).$ (1)
Load Balancing Loss. To encourage load balancing among the experts, we incorporate a differentiable load balancing loss [25, 54] into each MoE layer, promoting a more balanced distribution of input tokens across the experts. For a batch $\mathcal{B}$ containing $T$ tokens, the auxiliary loss is calculated as a scaled dot product between the vectors $f$ and $P$,
@@
$f_{i}=\frac{1}{T}\sum_{\boldsymbol{x}\in\mathcal{B}}\boldsymbol{1}\{\text{argmax }(\boldsymbol{x})=i\},$ (3)
$P_{i}=\frac{1}{T}\sum_{\boldsymbol{x}\in\mathcal{B}}g_{i}(\boldsymbol{x}).$ (4)
@@
MoE Transformer. A widely adopted approach for applying sparse MoE to Transformer [41] is to replace the feed-forward networks (FFNs) in certain standard (non-MoE) Transformer blocks with multiple experts [14, 36]. Specifically, the experts in the MoE layer retain the same structure as the original FFN. The gating network receives the output from the preceding self-attention layer and routes the tokens to different experts.
3.2 ViMoE
@@
We introduce a ViMoE framework to facilitate our study on the application of MoE in vision tasks. Specifically, we choose the Vision Transformer (ViT) [9] backbone and replace the FFNs in the ViT blocks with MoE layers. We consider inheriting self-supervised pre-training weights instead of training from scratch [36], which reduces training costs while benefiting from advanced pre-trained feature representations. Since the experts in the MoE layers share the same structure as the FFNs, we replicate the pre-trained weights of the FFNs across each expert for initialization.
@@
Shared Expert. There is often some common sense or shared information across input tokens assigned to different experts. As a result, with a conventional routing strategy, multiple experts may acquire overlapping knowledge within their respective parameters. By designing the shared expert [46, 6] to focus on capturing and consolidating common information, other routed experts can specialize in learning unique knowledge, leading to a more parameter-efficient model composed of a greater number of specialized experts. Consequently, we introduce the shared expert into ViMoE to learn common knowledge from all data. In our implementation, we set up one shared expert with the same structure as the other experts, whose output is added to the output of the selected routed experts.
Routing Strategy. Sparse MoE models typically employ a token-based routing strategy [36, 33, 6], where the gating mechanism allocates each token to selected experts. However, it is worth considering whether this strategy is suitable for vision MoE. For image classification, the model is expected to predict class based on the overall features of the image. Therefore, routing at the image level (i.e., selecting experts for the entire image) [7, 29] aligns more closely with the objectives of image classification. In practice, we use the [CLS] token to represent the image as input to the gating network since it encapsulates the information from all image tokens and is used for classification predictions. As to semantic segmentation, employing image-level routing is inappropriate; the token-based routing strategy better meets the requirements of pixel-level classification. We adapt the routing strategy in ViMoE tailored to different vision tasks, reflecting our suggestion that routing strategies should be congruent with the task objectives.
@@ 4 Empirical Observations in Designing ViMoE
In this section, we commence our study with image classification and present empirical observations and insightful phenomena encountered during the design of ViMoE.
@@ 4.1 A Stability Strategy for Convenient Design
Figure 3: Training curves for various ViMoE configurations.
@@
Figure 4: Routing heatmap of the $l$-th MoE layer, where $l=1$ represents the deepest (last) layer and $l=12$ denotes the shallowest (first) layer. The $x$-axis is the expert ID, and the $y$-axis is the class ID from ImageNet-1K. The label order in each figure is adjusted for better readability. Darker colors indicate a higher proportion of images from the corresponding class routed to the expert.
Scanning the Number of MoE Layers. An essential consideration in designing ViMoE is determining how many MoE layers to include and where to place them within the ViT blocks. For simplicity, we begin our exploration with a sparse MoE configuration without shared experts. The most straightforward approach is to place the MoE layer in every ViT block or to select the last $L$ blocks where the gradient magnitudes are the largest. To explore reasonable configurations and seek guiding insights, we scan the number of MoE layers and evaluate the classification accuracy. ViMoE employs the DINOv2 [32] pre-trained ViT-S/14 and is fine-tuned for $200$ epochs on ImageNet-1K [8] (more implementation details are provided in Sec. 6.1). From Fig. 2 (a), it can be observed that regardless of the number of experts, whether $N=2$, $N=4$, or $N=8$, the accuracy consistently exhibits a trend of initially increasing and then decreasing, with this trend becoming more pronounced as $N$ increases. This phenomenon has also been mentioned in Daxberger et al. [7]. We hypothesize that introducing multiple experts too early in the shallow ViT blocks leads to optimization difficulties, and the gating network struggles to achieve precise routing due to limited information (a more detailed analyze is given in Fig. 4). This suggests a potential instability in the design of ViMoE. Simply adding MoE layers to all ViT blocks without careful consideration may not lead to optimal results. A scan over different values of $L$ is required to determine the most suitable number of layers, which inevitably increases the design cost.
Shared Expert for Stabilising ViMoE. As previously discussed, the shared expert learns and consolidates knowledge from all the data, making it more effective in capturing common information. We consider this structure effective in alleviating the challenges of gating decisions and the limitations of individual expert learning within the sparse structure. Therefore, we attempt to incorporate the shared expert into ViMoE to mitigate the potential instability in training MoE layers. In Fig. 2 (b) we present a comparison between models with and without shared experts, where each MoE layer contains one shared expert. Incorporating the shared expert allows ViMoE to achieve stable results, eliminating the need for an exhaustive search to determine the optimal number of layers $L$. Even the naive approach of adding MoE layers to all ViT blocks yields good accuracy, preventing performance degradation caused by inappropriate MoE configurations. Additionally, with the inclusion of the shared expert, ViMoE achieves a 0.4% improvement in accuracy (84.3% vs. 83.9%), and a 1.2% increase compared to the DINOv2 baseline (83.1%).
Convergence Advantage. Using $N=8$ and $L=12$ as an example, Fig. 3 shows the training curves with and without shared experts, along with the DINOv2 baseline for reference. It is evident that simply adding sparse MoE layers slows down convergence in the early training epochs, and the final performance is nearly indistinguishable from the baseline, supporting the hypothesis that an improper MoE setting can even hinder optimization. In contrast, when shared experts are introduced, training becomes more stable, convergence is faster, and accuracy improves significantly. It is worth mentioning that, with the introduction of shared experts, each MoE layer contains a total of 9 experts (1 shared expert and 8 routed experts), and the forward pass activates both the shared expert and one selected routed expert. To ensure a fairer comparison, we conducted an ablation study by selecting the top-2 experts from the 9 routed experts. On the one hand, selecting 2 out of 9 can be seen as a denser setup than selecting 1 out of 8, which partially mitigates the adverse effects of being overly sparse. On the other hand, even with the same number of experts and activated experts, shared experts still demonstrate the advantage of faster convergence and higher accuracy.
4.2 Investigating Efficiency from Stable Structure
After constructing the stable ViMoE, we further analyze Fig. 2 (b) and observe a saturation phenomenon in performance. Interestingly, the inflection points vary with the number of experts $N$. For $N=2$, $N=4$, and $N=8$, accuracy already surpasses 84.2% at $L=5$, $L=3$, and $L=2$, respectively. Adding more MoE layers beyond these counts does not lead to significant improvements. We attempt to explain these phenomena and propose strategies for designing a more efficient ViMoE.
Routing Heatmap. Taking $N=8$ as an example, we plot the routing heatmaps of several MoE layers in Fig. 4. These heatmaps illustrate the distribution of class samples across different experts, helping us observe whether the experts are capable of capturing distinctive information. It can be observed that for the MoE layers in the shallow ViT blocks (e.g., $l=12$), the gating network struggles to consistently route images of the same class to the same expert or effectively distinguish the classes each expert should focus on. This indicates that the experts fail to learn highly discriminative knowledge; instead, they are likely performing similar functions, indiscriminately extracting common features. We then focus on the layer where the accuracy plateau occurs for $N=8$, corresponding to $L=2$. It is evident that in the last two MoE layers, the gating network can effectively assign the appropriate expert to each class, and the multiple experts can specialize in handling the corresponding data. Therefore, we conclude that the deep layers are where MoE truly achieves its divide-and-conquer objective, with different experts specializing in handling class-specific content. This observation validates the empirical approach of placing MoE layers in the last few ViT blocks [43, 29] as a reasonable strategy. In contrast, MoE struggles to demonstrate its advantages in the shallow ViT blocks, as the use of multiple experts seems unnecessary for capturing basic visual features. The sparse structure may instead introduce optimization difficulties, making the original dense FFN structure a simpler and more suitable choice.